# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v5

@@ Abstract
Swarms of autonomous interactive drones, with the support of recharging technology, can provide compelling sensing capabilities in Smart Cities, such as traffic monitoring and disaster response. This paper aims to deliver a novel coordination solution for the cost-effective navigation, sensing, and recharging of drones. Existing approaches, such as deep reinforcement learning (DRL), offer long-term adaptability, but lack energy efficiency, resilience, and flexibility in dynamic environments. Therefore, this paper proposes a novel approach where each drone independently determines its flying direction and recharging place using DRL, while adapting navigation and sensing through distributed optimization, which improves energy-efficiency during sensing tasks. Furthermore, drones efficiently exchange information while retaining decision-making autonomy via a structured tree communication model. Extensive experimentation with datasets generated from realistic urban mobility underscores an outstanding performance of the proposed solution compared to state-of-the-art methods. Significant new insights show that long-term methods optimize scarce drone resource for traffic management, while the integration of short-term methods is crucial for advising on charging policies and maintaining battery safety.
@@
drones, UAV, navigation, sensing, charging, distributed optimization, deep reinforcement learning
@@ I Introduction
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities [1]. With their mobility, autonomy, and diverse sensors, drones have been widely used in several applications, such as reporting traffic congestion at early stage [2], and mapping natural disasters [3]. However, the inherent limitations in battery capacity of drones influence their spatio-temporal coverage. On the other hand, it requires a high cost to build and maintain a large-scale drone swarm together. Therefore, efficient resource management is essential for long operations of drones by ensuring optimal drone recycling and preventing battery depletion. Recently, scholars have introduced the recharging technologies and commercially wireless charging stations for long-term sensing [4].
Inspired by these studies, this work tackles the problem of coordinating swarms of drones for improved distributed navigation, sensing and recharging. The goal is to enhance sensing quality in a large-scale area using both drones and charging stations while minimizing the overall involved costs. This endeavor requires highly adaptive and optimized drone operations. Previous work [5] employs distributed optimization methods for efficient drones path planning and task assignment to enhance sensing quality. These methods enable drones to autonomously self-organize and self-assign their tasks collectively, while respecting privacy and autonomy of drones. However, these methods are reactive to the current state of the environment rather than proactive to address long-term operational challenges strategically. The navigation and sensing of drones are significantly influenced by two factors: their flying directions and future requirements for sensor data on the map. For example, drones equipped with insights about the expected increase in traffic flow in certain areas can proactively fly to those locations, even if vehicle density is currently low. This strategic decision can be more effective to allocate limited drone resources for tasks like vehicle detection, although in short-term the benefits are not evident. This underscores the "slower is faster" effect, highlighting the need for long-term planning to optimize overall effectiveness and efficiency in drone sensing operations.
Deep reinforcement learning (DRL) tackles the challenge of long-term optimization. It accounts for long-term and iterative actions through the Bellman equation, which considers the discounted cumulative reward of agents’ actions [6, 7]. Drones, as heterogeneous agents with distinct behavior patterns and action spaces, can benefit from the powerful deep neural networks employed in DRL, enabling an effective handling of sophisticated state space and time-varying environments. Nevertheless, as the number of drones and time periods increase, the curse of dimensionality constrains efficient sensing. For instance, considering a scenario where drones are tasked with learning optimal path planning over a broad range of time scales using DRL. In this setting, the high number of epochs may inadvertently compel the drones to take unnecessary actions. As a result, drones could choose longer paths, leading to higher energy consumption. Furthermore, current methods overlook the need for drones to become more flexible and adaptive to real-time environmental changes while maintaining system resilience, against single points of failure. Addressing these limitations requires reducing actions in a temporal scale as well as effective autonomous mission planning distributively, but no existing approaches incorporate these novelties.
@@
Therefore, in this paper, we propose a new approach for the coordination of swarms of drones, Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a novel multi-agent DRL algorithm that builds upon on a collective learning approach  [8, 9]. DO-RL relies on a learning model to determine the drones’ strategic navigation and flying directions for sensing, which also determines the recharging stations, while leaving the self-planning and selection of specific sensing operations to distributed optimization, i.e., collective learning. This approach significantly reduces the number of decisions in a temporal scale, avoiding unnecessary energy consumption. Then, we design a structured tree communication model within which each agent interacts with its children and parent in a bottom-up and top-down fashion, obtains the aggregated observation to calculate immediate reward in DRL, and independently decides its navigation and sensing. This ensures scalability (support of a large number of agents), efficiency (low communication and computational cost), decentralization and resilience [10, 11, 12]. Furthermore, a set of auxiliary approaches is introduced, including a plan generation strategy that provides high-quality and diverse options for drones navigation and sensing, a plan selection, a periodic state update and a centralized training and decentralized execution framework. To evaluate the effectiveness of DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
@@
The contributions of this paper are outlined as follows: (i) The first attempt to coordinate navigation, sensing, and recharging by a swarm of autonomous and self-planning drones in a distributed and cost-effective way. (ii) A novel approach, DO-RL, to address this problem using both multi-agent DRL and distributed optimization. (iii) The applicability of a designed structured tree communication model for efficient planning and information sharing among agents to ensure decentralization, resilience, and flexibility. (iv) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (v) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (vi) An open-source implementation of the proposed approach and an open dataset [13] containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
@@ II Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as a combinatorial optimization problem, resembling the Traveling Salesmen Problem [5]. This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The digraph-based methods, such as greedy approach, [21], ant colony optimization [22], and genetic algorithm [23], are introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, inspection delay, events detection rate and the cost of flying trajectories. However, these works lack comprehensive evaluation in complex sensing scenarios, such as varying time periods, charging stations, and different sensing distributions. Moreover, these centralized methods suffer from a risk of single points of failure [24], wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms.
@@
Distributed optimization. Several distributed task assignment algorithms have been proposed to address the challenges of multi-UAV systems in complex sensing scenarios. One such algorithm is the robust decentralized task assignment (RDTA) [14] that enhances robustness and reduces communication costs by employing decentralized planning for drones. Another algorithm is the consensus-based bundle algorithm (CBBA) [15], which combines market-based mechanisms and situation awareness to converge and avoid task conflicts among drones. Furthermore, a planning-based coordination method using collective learning, named Economic Planning and Optimized Selections (EPOS), enables drones to optimize their navigation and sensing options under battery constraints [16, 25]. However, long-term sensing scenarios pose challenges. Drones tend to prioritize immediate maximum returns in the short term, often overlooking the need for strategic decision-making for decentralized navigation, sensing and recharging.
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective when addressing the distributed sensing problem with drones using complex optimization objectives, limited energy consumption, and a large number of heterogeneous agents. Afrin et al. [26] designed a DRL-based model for edge servers to dynamically allocate tasks to robots (e.g., drones), aimed at enhancing system resilience in smart farming environments. Ding et al. [17] integrated the mobile crowdsensing (i.e., human equipped with mobile sensing devices such as smartphones) into UAV sensing, which overcomes the limitations of both human mobility and drone battery capacity. Omoniwa et al. [18] presented a decentralized approach that leverages DRL and shared information from neighbors to improve the energy efficiency of drones in providing wireless connectivity to ground users in dynamic environments. Nevertheless, these approaches lack privacy and autonomy for drones in self-assignment tasks, reducing flexibility, adaptability to the real-time environmental changes and increasing vulnerability to single points of failure. Moreover, they are confined to a small number of drones and time periods due to the huge action space. In contrast, our approach empowers agents to determine overall flying directions every time period, leaving specific navigation and sensing tasks in each timeslot to planning and selection. This strategic allocation significantly filters out unnecessary actions and reduces the corresponding energy consumption.
Recharging assignment. Ealier work [27, 28] has studied the recharging scheduling problem: drones efficiently collect sensor data while charging from stations. However, scalability becomes a challenge since the algorithms used in these works only deal with a small number of drones. Zhao et al. [19] proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces, specifically for detecting available parking spots. Their solution incorporates the task selection and scheduling using DRL, and adaptive charging scheduling. However, this approach is centralized and tailored exclusively to the scenario of open parking spaces.
Transportation application. Despite the potential of drones in improving traffic monitoring accuracy, safety, and cost savings, there is limited research on task assignment in this particular scenario. Elloumi et al[29] proposed a road traffic monitoring system that generates adaptive drones trajectories by extracting information about vehicles. Huang et al. [30] developed the pre-flight conflict detection and resolution method, where service providers of drones with their own interest can coordinate to address the path planning problem. Samir et al. [20] leveraged a DRL algorithm to optimize the trajectories of UAVs, with the objective of minimizing the age of information of collected data. While these approaches demonstrate high performance in the detection of traffic vehicles, they do not explicitly address the challenges of decentralized sensing and recharging. Furthermore, they assume prior knowledge of the traffic flow of the vehicles, which may not be readily available in practical scenarios.
In summary, this paper bridges a gap in previous approaches. It empowers drones to make strategic choices in flying directions and recharging locations by predicting traffic flow, while enabling them to autonomously optimize their sensing operations, ensuring decentralization, resilience, and flexibility.
TABLE II: Notations.
Notation
Explanation
$u,U,\mathcal{U}$
Index of a drone; total number of drones; set of drones
$m,M,\mathcal{M}$
Index of a charging station; total number of charging stations; set of charging stations
$n,N,\mathcal{N}$
Index of a grid cell; total number of grid cells; set of grid cells
$s,S,\mathcal{S}$
Index of a timeslot; total number of timeslots in a period; set of timeslots
$t,T,\mathcal{T}$
Index of a period; total number of periods; set of periods
$a^{u}$
The action or flying direction taken by $u$
$P^{u},p^{u}_{n,s}$
The plan of $u$; value of $P^{u}$ at cell $n$ and timeslot $s$
$P^{-u},p^{-u}_{n,s}$
The observed plans of drones by $u$ excluding $P^{u}$; value of $P^{-u}$ at cell $n$ and timeslot $s$
$p_{n,s}$
The aggregated plans of all drones at cell $n$ and timeslot $s$
$V_{n,s}$
Required sensing value at cell $n$ and timeslot $s$
$v_{n,s}$
Collected sensing value at cell $n$ and timeslot $s$
$\mathcal{R},r_{n,s}$
Target; target value at cell $n$ and timeslot $s$
$l,L$
Index of a plan in $G^{a^{u}}$; total number of plans in $G^{a^{u}}$
$C^{\mathsf{f}}(u),t^{\mathsf{f}}$
Flying power consumption of $u$; flying time
$C^{\mathsf{h}}(u),t^{\mathsf{h}}_{u}$
Hovering power consumption of $u$; hovering time
$c^{u},e^{u}$
Current location of $u$; current energy consumption of $u$
$K(a^{u})$
Indexes of cells within the searching range along flying direction $a^{u}$
$J(a^{u})$
Indexes of visited cells within $K(a^{u})$
$\beta$
Behavior of an agent in planning optimization
$o^{u}_{t},r^{u}_{t}$
Observation of $u$ at period $t$; reward of $u$ at the period $t$
$k,H$
Index of a sampled transition; number of sampled transitions
$\gamma,\epsilon$
Discount factor; clip interval hyperparameter
$Q(\cdot),\theta^{Q}$
Critic network; parameter of critic network
$\pi(\cdot),\theta^{\pi}$
Actor network; parameter of actor network
$\omega_{n,s},\overline{v}$
Prediction coefficient to calculate $\hat{V}_{n,s}$; threshold to remove the regions with low sensing requirements
$I,\mathcal{X},A$
Total number of iterations in EPOS; size of state space; number of actions in DRL
$\mathcal{E},W$
Number of episodes; number of neutrals per layer in DNN
@@ III System Model
Table II illustrates the list of mathematical notations used in this paper.
@@ III-A Scenario and Assumptions
Matrix of plans. To explain the short-term navigation and sensing of drones over the cells and timeslots in a period $t$, $t\in\mathcal{T}$, the plan of a drone $u$ that travels through the direction of $a^{u}$ is introduced, denoted by $P^{u}(a^{u},t)$. It represents the specific navigation and sensing details, including the visited cells and corresponding energy consumption. The plan $P^{u}(a^{u},t)$ is encoded by a matrix of size $N\times S$, with each element represented as $p^{u}_{n,s}(a^{u},t)\in\{0,1\}$. Here, $p^{u}_{n,s}(a^{u},t)=1$ denotes that the drone $u$ hovers and collects all required data at cell $n$ at timeslot $s$, whereas for $p^{u}_{n,s}(a^{u},t)=0$ the drone does not hover at that cell at that time. Moreover, $P^{-u}(a^{u},t)=\{p^{-u}_{n,s}(a^{u},t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$ denotes the observed plans by drone $u$ in time period $t$, indicating that it incorporates and sums the plans of all other drones excluding its own. We formulate the aggregated plans of all drones observed by $u$ at cell $n$ and timeslot $s$ as follows:
@@
Assumptions. To simplify the scenario, this paper makes the following assumptions: (i) Each drone is programmed to fly at a unique altitude while moving between cells to prevent collisions [32]. We also fix the ground speed of drones and the wind speed for simply calculating power consumption. (ii) Each time period concludes a flying period and a charging period: the time that drones perform sensing, and the time for charging. All drones finish charging before the next time period begins. We assume that each charging period is of equal duration and provides sufficient time for the drones to be fully charged. (iii) Each charging station is adequately equipped with charging capacity, enabling multiple drones to charge simultaneously without the need for queuing. This arrangement prevents any delays in the charging process.

### v6

@@ Abstract
Swarms of autonomous interactive drones can provide compelling sensing capabilities in Smart City applications, such as traffic monitoring. This paper focuses on the task assignment problem for large-scale spatio-temporal sensing by a drone swarm. However, existing approaches have distinct challenges: evolutionary distributed optimization, such as collective learning, lacks long-term adaptability in dynamic environments, while deep reinforcement learning (DRL) is limited to scale effectively due to the curse of dimensionality. Therefore, this paper proposes a novel synergetic optimization approach by integrating long-term DRL and short-term collective learning. Through this approach, each drone independently and proactively determines its flying direction and recharging location using DRL, while evolving their navigation and sensing policies through collective learning based on a structured tree communication model. Extensive experiments with datasets generated from realistic urban mobility demonstrate an outstanding performance of the proposed solution in complex scenarios. New insights show that this approach provides a win-win synthesis of short-term and long-term strategies for drone-based traffic monitoring, with short-term methods addressing training complexity and energy management, while long-term methods preserving high sensing performance.
@@
drones, UAV, navigation, sensing, recharging, distributed optimization, deep reinforcement learning
@@ I Introduction
Unmanned Aerial Vehicles (UAVs), referred to as drones, can organize themselves into swarms, fostering collaboration and efficiency in sensor data collection within Smart Cities [1]. With their mobility, autonomy, and diverse sensors, drones have been widely used in several applications, such as monitoring traffic congestion [2] and mapping natural disasters [3]. Despite these advantages, drones are constrained by limited battery capacity, which impacts their spatio-temporal coverage. This limitation necessitates sophisticated autonomous control systems and precise task assignment within drone swarms. This involves the assignment of different sensing tasks to each drone while meeting the sensing requirements, drone capabilities and constraints [4]. Recently, scholars have solved the NP-hard combinatorial optimization problem of task assignment for spatio-temporal sensing, using evolutionary algorithms ranging from genetic algorithms to swarm intelligence methods [4, 5]. To enhance robustness and minimize the effects of individual drone failures, distributed optimization approaches have been introduced. These approaches allow drones to autonomously self-organize and self-assign tasks in a decentralized manner, preserving the autonomy of individual drones. However, these distributed approaches are reactive to the current state of the environment rather than proactive to address long-term operational challenges strategically.
Strategic foresight in navigation and sensing is crucial for optimizing drone operations in dynamic environments, significantly influenced by flying directions and future sensor data requirements. For example, drones aware of an expected increase in traffic can proactively fly to those areas, improving tasks such as vehicle detection, even if the short-term benefits are not evident. This "slower is faster" effect emphasizes the importance of long-term planning for overall effectiveness in drone sensing. To address long-term optimization, deep reinforcement learning (DRL) has been introduced, utilizing the Bellman equation to consider cumulative rewards [6]. DRL based on deep neural networks support drones, as heterogeneous agents, to manage complex state spaces and adapt to changing environments.
However, traditional DRL approaches cannot scale well with a large number of agents and extended time spans due to the curse of dimensionality. With more agents, coordinating them becomes exponentially more complex, leading to computational challenges. Furthermore, DRL methods need to account of environmental uncertainties and the long-term consequences of agents’ actions, which become increasingly complex as the time horizon extends. For example, drones often make flight path decisions over short intervals, such as seconds or minutes, but face significant challenges when required to adapt their policies for a long duration (e.g., a full day). Meanwhile, the extended operation is longer than the maximum flight time of drones, enabling them to return to a base station for recharging and depart from that place for the subsequent task. The selection of these recharging places constraints the drones’ flight ranges, which in turn affects their sensing performance.
@@
Therefore, in this paper, we address a novel task assignment problem for large-scale spatio-temporal sensing, focusing on the navigation (determining flight path), sensing (scheduling data collection) and recharging (choosing charging stations as destinations) of drone swarms in changing environments. To jointly manage these tasks, we propose a new synergetic optimization approach to coordinate drone swarms through Distributed Optimization and (deep) Reinforcement Learning (DO-RL), which involves a multi-agent DRL algorithm that builds upon on an iterative collective learning approach [7, 8]. It allows drones to determine long-term flying directions from departures to destinations (i.e., charging stations) via DRL, while autonomously evolving their short-term navigation and sensing policies based on shared plans from other drones through collective learning. The distributed sensing results generated during operations serve as rewards and observations for further DRL training. These results are aggregated and shared using a structured tree communication model, where each drone interacts with its parent and child nodes, making independent decisions. This ensures scalability (supporting a large number of agents), efficiency (low communication and computational cost), decentralization and resilience [9, 10, 11]. By employing this collaborative approach, collective learning reduces the frequency of decision-making and mitigates inefficient actions, effectively compensating for the limitations of DRL while retaining its strengths in long-term planning and optimization. In addition, a set of auxiliary approaches is introduced, including plan generation, plan selection, periodic state updates and a centralized training and decentralized execution module. To validate DO-RL, this paper conducts experimental evaluations with real-world data and emulated traffic monitoring scenarios, comparing it against state-of-the-art baseline methods.
@@
The contributions of this paper are outlined as follows: (i) The first study of task assignment for large-scale, dynamic spatio-temporal sensing by a swarm of drones, concentrating on optimizing the flight paths, data collection strategies, and recharging. (ii) A novel synergetic optimization approach, DO-RL, that tackles this problem by integrating both multi-agent DRL and distributed optimization based on a structured tree communication model. (iii) A testbed for extensive experimentation with realistic traffic patterns and real-world transport networks. (iv) The validation for the superior performance of the proposed approach over three state-of–the-art methods. (v) An open-source implementation of the proposed approach and an open dataset [12] containing all plans of the studied scenario. They can be used as benchmarks to encourage further research on this problem.
@@ II Related Work
Sensing task assignment. The UAV task assignment problem for spatio-temporal sensing has been traditionally defined as the Traveling Salesmen Problem [1, 20]. This involves finding the optimal assignment of tasks to drones at specific locations, while considering constraints such as task urgency, time scheduling, and flying costs. The evolutionary optimization methods, such as particle swarm optimization [21], and genetic algorithm [22], have been introduced to formulate the problem of reaching optimal sensing efficiency in terms of coverage, events detection rate and the cost of flight paths. They rely on adaptive exploration of the solution space to optimize task allocation based on shared information and local interactions. However, these methods are centralized, wherein the failure of a central control station can lead to several system disruptions without autonomous recovery mechanisms [23].
@@
Distributed optimization. Several distributed task assignment algorithms have been proposed. One such algorithm is the robust decentralized task assignment (RDTA) [13] that enhances robustness and reduces communication costs by employing decentralized planning for drones. Another one is the consensus-based bundle algorithm (CBBA) [14], which combines market-based mechanisms and situation awareness to converge and avoid task conflicts among drones. Moreover, a planning-based coordination method using collective learning, named Economic Planning and Optimized Selections (EPOS), enables drones to optimize their navigation and sensing options under battery constraints [15, 24]. However, long-term sensing scenarios pose challenges. Drones tend to prioritize immediate maximum returns in the short term, often neglecting the need for strategic decision-making for overall benefits.
DRL-based drone sensing. The deep reinforcement learning (DRL) algorithm proves to be effective for addressing distributed sensing with drones, managing complex optimization goals, energy constraints, and numerous heterogeneous agents. Ding et al. [16] integrated mobile crowdsensing (i.e., human equipped with mobile sensing devices) into UAV sensing, overcoming the limitations of both human mobility and drone battery life. Omoniwa et al. [25] presented a decentralized DRL approach that improves drone energy efficiency in providing wireless connectivity by sharing information with neighbouring drones. Zhao et al. [17] proposed a comprehensive solution for sustainable urban-scale sensing in open parking spaces. Their solution incorporates the task selection and scheduling using DRL, and adaptive recharging assignment. However, these approaches lack autonomy for drones in self-assignment tasks, reducing adaptability to the real-time environmental changes and increasing vulnerability to agent failures [23]. Additionally, they are confined to a small-scale task assignment problem due to the training complexity.
Evolutionary algorithm and DRL. In this optimization process, the related synergetic optimization algorithms integrate the full learning and optimization capabilities of evolutionary algorithms and DRL. They independently solve sub-problems to generate partial solutions, which are then combined to form a complete solution [26]. Earlier work [27, 28] has proposed optimizing team policies using evolutionary algorithms with team rewards, while simultaneously optimizing individual policies through DRL with agent rewards, thus bypassing the non-stationary problem in multi-agent DRL. Li et al. [18] introduce a hybrid framework using shared observation encoders and independent linear policy representations to enhance collaboration on complex task, ensuring stable evolution through agent-level crossover and mutation. While benchmarks, such as MuJoCo [29] and Starcraft II [30], are useful for evaluation of these approaches, these benchmarks may struggle to adapt to the continuously evolving and unanticipated dynamics typical of real-world scenarios, such as fluctuating traffic patterns in drone-based monitoring systems.
Transportation application. In the application of transportation system, drones rely on accurate vehicle observation to detect early signs of traffic congestion. This enables traffic operators to apply mitigation measures, reducing the carbon footprint in one of the world’s highest-emission sectors. Elloumi et al[31] proposed a road traffic monitoring system that generates adaptive drones trajectories by extracting information about vehicles. Samir et al. [19] leveraged a DRL algorithm to optimize the trajectories of UAVs, with the objective of minimizing the age of information of collected data. While these approaches demonstrate high performance in the detection of traffic vehicles, scalability becomes a challenge since they only deal with a small number of drones within a short duration.
In summary, previous approaches have either neglected the long-term advantages of strategic sensing, or failed to effectively adapt to failures and scale without encountering centralized bottlenecks. To overcome these shortcomings, the proposed approach empowers drones to make informed and proactive decisions based on predicted traffic flow, while enabling them to independently adjust to changing conditions and unexpected events, ensuring both scalability and resilience without relying on centralized control.
@@ III System Model
In this section, we define key concepts of scenarios and then formulate the main performance metrics.
@@ III-A Scenario and Assumptions
Matrix of plans. To explain the short-term navigation and sensing of drones over the cells and timeslots in a period $t$, $t\in\mathcal{T}$, the plan of a drone $u$ that travels through the direction of $a^{u}$ is introduced, denoted by $P^{u}(a^{u},t)$. It represents the specific navigation and sensing details, including the visited cells and corresponding energy consumption. The plan $P^{u}(a^{u},t)$ is encoded by a matrix of size $N\times S$, with each element represented as $p^{u}_{n,s}(a^{u},t)\in\{0,1\}$. Here, $p^{u}_{n,s}(a^{u},t)=1$ denotes that the drone $u$ hovers and collects all required data at cell $n$ at timeslot $s$, whereas for $p^{u}_{n,s}(a^{u},t)=0$ the drone does not hover at that cell at that time. Moreover, $P^{-u}(a^{u},t)=\{p^{-u}_{n,s}(a^{u},t)|_{n\in\mathcal{N},s\in\mathcal{S}}\}$ denotes the observed plan by drone $u$ in time period $t$, indicating that it incorporates and sums the plans of all other drones excluding its own. We formulate the aggregated plans of all drones observed by $u$ at cell $n$ and timeslot $s$ as follows:
@@
Assumptions. To simplify the scenario, this paper makes the following assumptions: (i) Each drone is programmed to fly at a distinct altitude during its movement between cells to prevent mid-air collisions risks [33]. While previous work [34, 35] has explored a safe and cost-effective approach to a more realistic collision avoidance, this is not the primary focus of this paper. (ii) Each time period concludes a flying period and a charging period: the time that drones perform sensing, and the time for charging. All drones finish charging before the next time period begins. We assume that each charging period is of equal duration and provides sufficient time for the drones to be fully charged. (iii) Each charging station is adequately equipped with charging capacity, enabling multiple drones to charge simultaneously without the need for queuing. This arrangement prevents any delays in the charging process.